2016 2017 2016 2017 Current assets $ 934 $ 984 Current liabilities $ 375 $ 398 Net fixed assets 3,897 4,576 Long-term debt 2,021 2,167 WESTON ENTERPRISES 2017 Income Statement Sales $ 11,770 Costs 5,710 Depreciation 1,060 Interest paid 180 a. What was owners' equity for 2016 and 2017? (Do not round intermediate calculations.) Owners' equity 2016 $ Owners' equity 2017 $ b. What was the change in net working capital for 2017? (Do not round intermediate calculations.) Change in NWC $ c-1 In 2017, the company purchased $1,845 in new fixed assets. How much in fixed assets did the company sell? (Do not round intermediate calculations.) Fixed assets sold $ c-2 In 2017, the company purchased $1,845 in new fixed assets. What was the cash flow from assets for the year? The tax rate is 35 percent. (Do not round intermediate calculations.) Cash flow from assets $ d-1 During 2017, the company raised $400 in new long-term debt. How much long-term debt must the company have paid off during the year? (Do not round intermediate calculations.) Debt retired $ d-2 During 2017, the company raised $400 in new long-term debt. What was the cash flow to creditors? (Do not round intermediate calculations.) Cash flow to creditors $
Answer #1
2016 2017
Current assets 984 984
Net fixed assets 3897 4576
Total assets 4881 5560
Current liabilities 375 378
Long term debt 2021 2167
Total debt 2396 2545
Equity (Plug in figure) 2485 3015
Total liabilities and Equity 4881 5560
Sales 11770
Costs 5710
Depreciation 1060
EBIT 5000
Interest paid 180
EBT 4820
Tax at 35% 1687
NI 3133
Dividends (3133-530) 2603
Addition to retained earnings (3015-2485) 530
ANSWERS:
a) Owners' equity 2485 3015
b) Change in NWC = (984-378)-(984-375) = -3
c-1) Fixed assets sold = (3897-1060+1845)-4576 = 106
c-2) Cash flow from assets = OCF-Net capital spending-Change in NWC
OCF = NI+Depreciation+Interest = 3133+180+1060 = 4373
Cash flow from assets = 4373-(1845-106)+3 = 2637
d-1) LT debt paid off = (2021+400)-2167 = 254
Cash flow to creditors = 254+180-400 = 34
Cash flow to stockholders 2603
CHECK:
Cash flow from assets = Cash flow to creditors + Cash flow to stockholders
2637 = 34 + 2603